OpenStudy (anonymous):

Simplify. sqrt108-3sqrt8-sqrt27

OpenStudy (anonymous):

\[108=36\times 3\] so \[\sqrt{108}=\sqrt{36\times 3}=\sqrt{36}\times \sqrt{3}=6\sqrt{3}\]

OpenStudy (anonymous):

\[8=4\times 2\] so \[-3\sqrt{8}=-3\times \sqrt{4}\times \sqrt{2}=-3\times 2\times \sqrt{2}=-6\sqrt{2}\]

OpenStudy (anonymous):

and finally \[27=9\times 3\] so ...

OpenStudy (anonymous):

-3√3-6√2!